OpenStudy (anonymous):

se factoring to solve the quadratic equation: 4x^2=36x-81 im so bad in math =(

jhonyy9 (jhonyy9):

4x2-36x+81=0 x_1,_2=(36+/- sqrt(1296-1296))/8 = 36/8 = 9/2 factoring will be (2x-9) squared than (2x-9)(2x-9)